# Insert-Only versus Insert-Delete in Dynamic Query Evaluation

### v2

@@ 3.2. Insert-Delete Setting
In the insert-delete setting, our approach can maintain arbitrary join queries, and in particular any cyclic join query, with update times that can be asymptotically lower than recomputation. We give an upper bound on the update time for a query $Q$ in terms of the fractional hypertree width of the multivariate extension $\widehat{Q}$ of $Q$ (Definition 2.3):
Theorem 3.4 ().
[…]
@@
Corollary 3.5 ().
[…]
Sec. 7.1 overviews our algorithm that meets the upper bound in Theorem 3.4. It involves intersection joins (Sec. 6). The following statements shed light on the relationship between a query $Q$ and its multivariate extension $\widehat{Q}$ (their proofs are deferred to Appendix E):
  • $Q$ is hierarchical if and only if its multivariate extension $\widehat{Q}$ is acyclic, which is equivalent to $\mathsf{w}(\widehat{Q})=1$.
@@
A close analysis of the lower bound proof for non-hierarchical queries in prior work (Berkholz et al., 2017) reveals that for any non-hierarchical query $Q$ and any $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-conjecture fails (Proposition E.5 in Appendix E). Following Theorem 3.4 and the fact that $\mathsf{w}(\widehat{Q})=\frac{3}{2}$ for any Loomis-Whitney query $Q$, we conclude that both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{1/2})$ amortized update time, which is optimal, unless the OMv-conjecture fails.
[…]
@@
Theorem 3.6 ().
[…]
@@ 4. Preliminaries
Acyclic queries are reviewed in Appendix A.
@@
A union of queries is acyclic (hierarchical, Loomis-Whitney) if each of its components is acyclic (hierarchical, Loomis-Whitney). Example A.1 showcases the query classes mentioned above.
@@
Given a query $Q$ and a database $\mathcal{D}$, an enumeration procedure for $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) outputs the elements in $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) one by one in some order and without repetition. The enumeration delay of the procedure is the maximum of three times: the time between the start of the enumeration process and the output of the first element, the time between outputting any two consecutive elements, and the time between outputting the last element and the end of the enumeration process (Durand and Grandjean, 2007). The enumeration delay is constant if it does not depend on $|\mathcal{D}|$.
@@
The AGM-bound and fractional hypertree width of join queries are defined as usual. Their definitions are deferred to Appendix A. If a query is a union of join queries, its AGM-bound (or fractional hypertree width) is the maximum AGM-bound (or fractional hypertree width) over its components.
@@ 5. IVM: Insert-Only Setting
In this section, we give an overview of our algorithm for $\text{IVM}^{+}[Q]$ that meets the upper bound of Theorem 3.1. We leave the details and proofs (including the extension to $\text{IVM}^{+}_{\delta}[Q]$) to Appendix B. We start with the following lemma: (Recall notation from Sec. 4.)
Lemma 5.1 ().
@@
Note that the total computation time above is the same as the AGM bound (Atserias et al., 2013) of $Q$ over the final database $\mathcal{D}^{(N)}$. In particular, even in the static setting where $\mathcal{D}^{(N)}$ is given upfront, the output size $|Q(\mathcal{D}^{(N)})|$ could be as large as $\text{\sf AGM}(Q,\mathcal{D}^{(N)})$ in the worst-case. Moreover, worst-case optimal join algorithms cannot beat this runtime (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). The above lemma is proved based on the query decomposition lemma (Ngo et al., 2014; Ngo, 2018), which says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$. Appendix B gives a proof of the query decomposition lemma and an example.
@@
$\min(|\sigma_{B=b}S^{(\tau)}(B,C)|,|\sigma_{A=a}T^{(\tau)}(A,C)|)\leq$
(7) $\sqrt{|\sigma_{B=b}S^{(\tau)}(B,C)|\cdot|\sigma_{A=a}T^{(\tau)}(A,C)|}$
@@
The query decomposition lemma (specifically Eq. (39) in Example B.2) says that the sum of the quantity in Eq. (8) over all $(a,b)\in R^{(N)}$ is bounded by $\sqrt{|R^{(N)}|\cdot|S^{(N)}|\cdot|T^{(N)}|}\leq N^{3/2}$. Hence, all inserts into $R$ take time $\mathcal{O}(N^{3/2})$.
@@
In the above example, we only needed to do the calibration bottom-up from the leaf $Q_{1}$ to the root $Q_{2}^{\prime}$. We show in Appendix B.4 that this is not sufficient for $\text{IVM}^{+}_{\delta}[Q]$. Instead, we also need to do an extra calibration top-down. For example, suppose we have an insert of a tuple $(a,b)$ into $R$ above. In order to enumerate the new output tuples corresponding to $(a,b)$, we have to start our enumeration from $Q_{1}$ as the root. For that purpose, $Q_{1}$ also needs to be calibrated with $Q_{2}^{\prime}$. See Example B.3.
@@ 6. Technical Tools: Intersection Joins
Example 6.3 ().
Consider the following intersection query, which is the time extension of $Q_{\triangle}$ from Eq. (1):
@@
Let $\overline{Q}$ be the time extension of a query $Q$. Prior work reduces the evaluation of $\overline{Q}$ to the evaluation of $\widehat{Q}$, where $\widehat{Q}$ is the multivariate extension of $Q$ (Eq. (3)), whose variables are all point variables (Khamis et al., 2022). Moreover, it is shown in (Khamis et al., 2022) that this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown by a backward reduction from (the static evaluation of) each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail in Appendix C and briefly summarize the forward reduction below. We rely on both reductions to prove upper and lower bounds for IVM in the insert-delete setting (Sec. 3.2).
@@
Suppose we want to evaluate $\overline{Q}$ on a given database instance $\overline{\mathcal{D}}$. We construct a segment tree, $\mathcal{T}_{N}$, over the interval variable $[Z]$, where the parameter $N$ is roughly the maximum number of different $Z$-values. A segment tree $\mathcal{T}_{N}$ is a balanced binary tree with $N$ leaves. Each node corresponds to an interval. The root corresponds to the interval $[1,N]$, its left and right child correspond to the intervals $[1,N/2]$ and $[N/2+1,N]$ respectively, and so on. Each node can be represented by a bitstring of length at most $\log_{2}N$. Figure 2 in Appendix C depicts the segment tree $\mathcal{T}_{8}$. Each interval $[z]$ can be broken down into at most $\mathcal{O}(\log N)$ nodes in $\mathcal{T}_{N}$. We refer to these nodes as the canonical partition of $[z]$, and denote them by $\text{\sf CP}_{N}([z])$. For example, the canonical partitions of $[2,8]$ and $[2,5]$ in the segment tree $\mathcal{T}_{8}$ are:
@@
The canonical partition of the intersection of some intervals is roughly the intersection of their canonical partitions (Lemma C.1). Two (or more) nodes in the segment tree correspond to overlapping intervals if and only if the corresponding bitstrings have a prefix-suffix relationship (Lemma C.1). For example, the nodes labeled by $1$ and $100$ in Figure 2 correspond to two overlapping intervals, namely $[5,8]$ and $[5,5]$, because the string $1$ is a prefix of $100$. Therefore, testing intersections can be reduced to testing whether some bitstrings form a chain of prefixes. Armed with this idea, we convert $\overline{\mathcal{D}}$ into a database instance $\widehat{\mathcal{D}}$ for $\widehat{Q}$, called the canonical partition of $\overline{\mathcal{D}}$ and denoted by $\text{\sf CP}_{N}(\overline{\mathcal{D}})$. There is a mapping between the outputs $\overline{Q}(\overline{\mathcal{D}})$ and $\widehat{Q}(\widehat{\mathcal{D}})$ (Theorem C.5).
Example 6.4 ().
Suppose we want to reduce $\overline{Q}_{\triangle}$ from Eq. (11) to $\widehat{Q}_{\triangle}$ from Example 2.4. In particular, we are given the database instance $\overline{\mathcal{D}}$ from Figure 1(d) and want to compute the corresponding instance $\widehat{\mathcal{D}}$ for $\widehat{Q}_{\triangle}$. We show how to construct $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ for $\widehat{Q}_{123}$ in Eq. (4). The remaining $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$ are similar. The query $\widehat{Q}_{123}$ is meant to test whether $\overline{R},\overline{S},\overline{T}$ contain three intervals $[z^{R}],[z^{S}],[z^{T}]$ whose canonical partitions contain bitstrings $z^{R},z^{S},z^{T}$ where $z^{R}$ is a prefix of $z^{S}$ which is a prefix of $z^{T}$. This can only happen if there are bitstrings $z_{2},z_{3}$ where $z^{S}=z^{R}\circ z^{2}$ and $z^{T}=z^{S}\circ z^{3}$, where $\circ$ denotes string concatenation. We define:
@@
$\widehat{T}_{123}$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge$
$\quad\quad\quad\quad(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
Let $\varepsilon$ denote the empty string. Applying the above to $\overline{\mathcal{D}}$ from Figure 1(d), $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ respectively contain the tuples: $(\varepsilon,a_{1},b_{1})$, $(\varepsilon,01,b_{1},c_{1})$, $(\varepsilon,01,\varepsilon,a_{1},c_{1})$, which join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$. This is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. See Example C.2 for more details.
@@ 7. IVM: Insert-Delete Setting
In this section, we give a brief overview of how we obtain our upper and lower bounds for IVM in the insert-delete setting from Theorems 3.4 and 3.6, respectively. Details are deferred to Appendix D.

### v3

@@ 3.2. Insert-Delete Setting
In the insert-delete setting, our approach can maintain arbitrary join queries, and in particular any cyclic join query, with update times that can be asymptotically lower than recomputation. In particular, we give an upper bound on the update time for a query $Q$ in terms of the fractional hypertree width of the multivariate extension $\widehat{Q}$ of $Q$ (Definition 2.3):
Theorem 3.4.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(N^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$ and $N$ is the number of single-tuple updates.
@@
Corollary 3.5.
For any query $Q$, both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}^{(\tau)}|^{\mathsf{w}(\widehat{Q})-1})$ amortized update time, where $\widehat{Q}$ is the multivariate extension of $Q$, and $|\mathcal{D}^{(\tau)}|$ is the current database size at update time $\tau$.
Sec. 7.1 overviews our algorithm that meets the upper bound in Theorem 3.4. It involves intersection joins (Sec. 6). The following statements shed light on the relationship between a query $Q$ and its multivariate extension $\widehat{Q}$ (their proofs are deferred to Appendix F):
  • $Q$ is hierarchical if and only if its multivariate extension $\widehat{Q}$ is acyclic, or equivalently $\mathsf{w}(\widehat{Q})=1$.
@@
A close analysis of the lower bound proof for non-hierarchical queries in prior work (Berkholz et al., 2017) reveals that for any non-hierarchical query $Q$ and any $\gamma>0$, there is no algorithm that solves $\text{IVM}^{\pm}[Q]$ or $\text{IVM}^{\pm}_{\delta}[Q]$ with $\widetilde{\mathcal{O}}(N^{1/2-\gamma})$ amortized update time, unless the OMv-conjecture fails (Proposition F.5 in Appendix F). Following Corollary 3.5 and the fact that $\mathsf{w}(\widehat{Q})=\frac{3}{2}$ for any Loomis-Whitney query $Q$, we conclude that both $\text{IVM}^{\pm}[Q]$ and $\text{IVM}^{\pm}_{\delta}[Q]$ can be solved with $\widetilde{\mathcal{O}}(|\mathcal{D}|^{1/2})$ amortized update time, which is optimal, unless the OMv conjecture fails.
First-order IVM, i.e., delta queries, and even higher-order IVM, i.e., delta queries with materialized views, cannot achieve the update time of our approach. It was already discussed in prior work that for the triangle join query, both IVM approaches need linear update time per single-tuple update (Kara et al., 2019, 2020a). See Appendix A. The IVMϵ approach resorts to a heavy/light partitioning argument that is tailored to the triangle query and does not generalize to other cyclic queries (Kara et al., 2019, 2020a). Instead, our approach solves this problem more systematically and for any query $Q$ by translating the temporal dimension (the tuple lifespan as defined by its insert and possible delete) into spatial attributes (the multivariate encoding of the tuple lifespan), taking a tree decomposition of the multivariate extension $\hat{Q}$ of $Q$, and by materializing and maintaining the bags of this tree decomposition.
@@
Theorem 3.6.
Let $Q$ be a query and $\widehat{Q}_{\bm{\sigma}}$ any component of its multivariate extension. For any constant $\gamma>0$, $\text{IVM}^{\pm}_{\delta}[Q]$ cannot be solved with amortized update time $\widetilde{\mathcal{O}}(N^{\mathsf{\omega}(\widehat{Q}_{\bm{\sigma}})-1-\gamma})$.
@@ 4. Preliminaries
Acyclic queries are reviewed in Appendix B.
@@
A union of queries is acyclic (hierarchical, Loomis-Whitney) if each of its components is acyclic (hierarchical, Loomis-Whitney). Example B.1 showcases the query classes mentioned above.
@@
Given a query $Q$ and a database $\mathcal{D}$, an enumeration procedure for $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) outputs the elements in $Q(\mathcal{D})$ (or $\delta_{\tau}Q(\mathcal{D})$) one by one in some order and without repetition. The enumeration delay of the procedure is the maximum of three times: the time between the start of the enumeration process and the output of the first element, the time between outputting any two consecutive elements, and the time between outputting the last element and the end of enumeration (Durand and Grandjean, 2007). The enumeration delay is constant if it does not depend on $|\mathcal{D}|$.
@@
The AGM-bound and fractional hypertree width of join queries are defined as usual. Their definitions are deferred to Appendix B. If a query is a union of join queries, its AGM-bound (or fractional hypertree width) is the maximum AGM-bound (or fractional hypertree width) over its components.
@@ 5. IVM: Insert-Only Setting
In this section, we give an overview of our algorithm for $\text{IVM}^{+}[Q]$ that meets the upper bound of Theorem 3.1. We leave the details and proofs (including the extension to $\text{IVM}^{+}_{\delta}[Q]$) to Appendix C. We start with the following lemma: (Recall notation from Sec. 4.)
Lemma 5.1.
@@
Note that the total computation time above is the same as the AGM bound (Atserias et al., 2013) of $Q$ over the final database $\mathcal{D}^{(N)}$. In particular, even in the static setting where $\mathcal{D}^{(N)}$ is given upfront, the output size $|Q(\mathcal{D}^{(N)})|$ could be as large as $\text{\sf AGM}(Q,\mathcal{D}^{(N)})$ in the worst-case. Moreover, worst-case optimal join algorithms cannot beat this runtime (Ngo et al., 2018; Veldhuizen, 2014; Ngo, 2018). The above lemma is proved based on the query decomposition lemma (Ngo et al., 2014; Ngo, 2018), which says the following. Let $Q$ be a query and let $\bm{Y}\subseteq\text{\sf vars}(Q)$. Then, the AGM bound of $Q$ can be decomposed into a sum of AGM bounds of “residual” queries: one query $Q\Join\bm{y}$ for each tuple $\bm{y}$ over the variables $\bm{Y}$. Appendix C gives a proof of the query decomposition lemma and an example.
@@
(7) $\min(|\sigma_{B=b}S^{(\tau)}(B,C)|,|\sigma_{A=a}T^{(\tau)}(A,C)|)\leq\sqrt{|\sigma_{B=b}S^{(\tau)}(B,C)|\cdot|\sigma_{A=a}T^{(\tau)}(A,C)|}$
@@
The query decomposition lemma (specifically Eq. (29) in Example C.2) says that the sum of the quantity in Eq. (8) over all $(a,b)\in R^{(N)}$ is bounded by $\sqrt{|R^{(N)}|\cdot|S^{(N)}|\cdot|T^{(N)}|}\leq N^{3/2}$. Hence, all inserts into $R$ take time $\mathcal{O}(N^{3/2})$.
@@
In the above example, we only needed to do the calibration bottom-up from the leaf $Q_{1}$ to the root $Q_{2}^{\prime}$. We show in Appendix C.4 that this is not sufficient for $\text{IVM}^{+}_{\delta}[Q]$. Instead, we also need to do an extra calibration top-down. For example, suppose we have an insert of a tuple $(a,b)$ into $R$ above. In order to enumerate the new output tuples corresponding to $(a,b)$, we have to start our enumeration from $Q_{1}$ as the root. For that purpose, $Q_{1}$ also needs to be calibrated with $Q_{2}^{\prime}$. See Example C.3.
@@ 6. Technical Tools: Intersection Joins
Example 6.3.
Consider this intersection query, which is the time extension of $Q_{\triangle}$ from Eq. (1):
@@
Let $\overline{Q}$ be the time extension of a query $Q$. Prior work reduces the evaluation of $\overline{Q}$ to the evaluation of $\widehat{Q}$, where $\widehat{Q}$ is the multivariate extension of $Q$ (Eq. (3)), whose variables are all point variables (Khamis et al., 2022). Moreover, it is shown in (Khamis et al., 2022) that this reduction is optimal (up to a $\widetilde{\mathcal{O}}(1)$ factor). In particular, the query $\overline{Q}$ is exactly as hard as the hardest component $\widehat{Q}_{\bm{\sigma}}$ in $\widehat{Q}$. This is shown by a backward reduction from (the static evaluation of) each component $\widehat{Q}_{\bm{\sigma}}$ to $\overline{Q}$. We explain both reductions in detail in Appendix D and briefly summarize the forward reduction below. We rely on both reductions to prove upper and lower bounds for IVM in the insert-delete setting (Sec. 3.2).
@@
Suppose we want to evaluate $\overline{Q}$ on a given database instance $\overline{\mathcal{D}}$. We construct a segment tree, $\mathcal{T}_{N}$, over the interval variable $[Z]$, where the parameter $N$ is roughly the maximum number of different $Z$-values. A segment tree $\mathcal{T}_{N}$ is a balanced binary tree with $N$ leaves. Each node corresponds to an interval. The root corresponds to the interval $[1,N]$, its left and right child correspond to the intervals $[1,N/2]$ and $[N/2+1,N]$ respectively, and so on. Each node can be represented by a bitstring of length at most $\log_{2}N$. Figure 10 in Appendix D depicts the segment tree $\mathcal{T}_{8}$. Each interval $[z]$ can be broken down into at most $\mathcal{O}(\log N)$ nodes in $\mathcal{T}_{N}$. We refer to these nodes as the canonical partition of $[z]$, and denote them by $\text{\sf CP}_{N}([z])$. For example, the canonical partitions of $[2,8]$ and $[2,5]$ in the segment tree $\mathcal{T}_{8}$ are:
@@
The canonical partition of the intersection of some intervals is roughly the intersection of their canonical partitions (Lemma D.1). Two (or more) nodes in the segment tree correspond to overlapping intervals if and only if the corresponding bitstrings have a prefix-suffix relationship (Lemma D.1). For example, the nodes labeled by $1$ and $100$ in Figure 10 correspond to two overlapping intervals, namely $[5,8]$ and $[5,5]$, because the string $1$ is a prefix of $100$. Therefore, testing intersections can be reduced to testing whether some bitstrings form a chain of prefixes. Armed with this idea, we convert $\overline{\mathcal{D}}$ into a database instance $\widehat{\mathcal{D}}$ for $\widehat{Q}$, called the canonical partition of $\overline{\mathcal{D}}$ and denoted by $\text{\sf CP}_{N}(\overline{\mathcal{D}})$. There is a mapping between the outputs $\overline{Q}(\overline{\mathcal{D}})$ and $\widehat{Q}(\widehat{\mathcal{D}})$ (Theorem D.5).
Example 6.4.
Suppose we want to reduce $\overline{Q}_{\triangle}$ from Eq. (11) to $\widehat{Q}_{\triangle}$ from Example 2.4. In particular, we are given the database instance $\overline{\mathcal{D}}$ from Figure 2(d) and want to compute the corresponding instance $\widehat{\mathcal{D}}$ for $\widehat{Q}_{\triangle}$. We show how to construct $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ for $\widehat{Q}_{123}$ in Eq. (4). The remaining $\widehat{Q}_{132},\ldots,\widehat{Q}_{321}$ are similar. The query $\widehat{Q}_{123}$ is meant to test whether $\overline{R},\overline{S},\overline{T}$ contain three intervals $[z^{R}],[z^{S}],[z^{T}]$ whose canonical partitions contain bitstrings $z^{R},z^{S},z^{T}$ where $z^{R}$ is a prefix of $z^{S}$ which is a prefix of $z^{T}$. This can only happen if there are bitstrings $z_{2},z_{3}$ where $z^{S}=z^{R}\circ z_{2}$ and $z^{T}=z^{S}\circ z_{3}$, where $\circ$ denotes string concatenation. We define:
@@
$\widehat{T}_{123}$ =$\stackrel{{\scriptstyle\text{def}}}{{=}}\{(z_{1},z_{2},z_{3},a,c)\mid\exists z:([z],a,c)\in\overline{T}\wedge(z_{1}\circ z_{2}\circ z_{3})\in\text{\sf CP}_{N}([z])\}$
Let $\varepsilon$ denote the empty string. Applying the above to $\overline{\mathcal{D}}$ from Figure 2(d), $\widehat{R}_{123},\widehat{S}_{123},\widehat{T}_{123}$ respectively contain the tuples: $(\varepsilon,a_{1},b_{1})$, $(\varepsilon,01,b_{1},c_{1})$, $(\varepsilon,01,\varepsilon,a_{1},c_{1})$, which join together producing the output tuple $(\varepsilon,01,\varepsilon,a_{1},b_{1},c_{1})$ of $\widehat{Q}_{123}$. This is a witness to the output tuple $([3,5],a_{1},b_{1},c_{1})$ of $\overline{Q}_{\triangle}$. See Example D.2 for more details.
@@ 7. IVM: Insert-Delete Setting
In this section, we give a brief overview of how we obtain our upper and lower bounds for IVM in the insert-delete setting from Theorems 3.4 and 3.6, respectively. Details are deferred to Appendix E.